The following selected transactions were taken from the books of Ripley Company for Year 1: On February 1, Year 1, borrowed $59,000 cash from the local bank. The note had a 5 percent interest rate and was due on June 1, Year 1. Cash sales for the year amounted to $230,000 plus sales tax at the rate of 6 percent. Ripley provides a 90-day warranty on the merchandise sold. The warranty expense is estimated to be 3 percent of sales. Paid the sales tax to the state sales tax agency on $205,000 of the sales. Paid the note due on June 1 and the related interest. On November 1, Year 1, borrowed $37,000 cash from the local bank. The note had a 7 percent interest rate and a one-year term to maturity. Paid $3,300 in warranty repairs. A customer has filed a lawsuit against Ripley for $90,000 for breach of contract. The company attorney does not believe the suit has merit. Required a. Answer the following questions: 1. What amount of cash did Ripley pay for interest during Year 1? 2. What amount of interest expense is reported on Ripley’s income statement for Year 1? 3. What is the amount of warranty expense for Year 1? b. Post the liabilities transactions to T-accounts and prepare the current liabilities section of the balance sheet at December 31, Year 1. c. Show the effect of these transactions on the financial statements using a horizontal statements model like the one shown next. Use + for increase, − for decrease, and NA for not affected. In the Cash Flow column, indicate whether the item is an operating activity (OA), investing activity (IA), or financing activity (FA) or not affected (NA).
Answer #1

a.

1 Cash paid for interest $           983
2 Interest expense $       1,415
3 Warranty expense $       6,900

Cash paid for interest = $59000 x 5% x 4/12 = $983.33

Interest expense = $983.33 + ($37000 x 7% x 2/12) = $983.33 + $431.67 = $1415

Warranty expense = 3% x $230000 = $6900

b.

Interest Payable Sales Tax Payable
432 6 13800 2
Bal. 432 4 12300
Bal. 1500
$37000 x 7% x 2/12 = $432 $230000 x 6% = $13800
$205000 x 6% = $12300
Warranty Payable Notes Payable
6900 3 59000 1
7 3300 5 59000
Bal. 3600 37000 6
$230000 x 3% = $6900 Bal. 37000

The lawsuit filed against Ripley by a customer for $90,000 for breach of contract should not be recorded since it is a contingent liability with probability of its occurrence being remote as the attorney believes the suit does not have merit.

RIPLEY COMPANY
Balance Sheet (Partial)
As of December 31, Year 1
Current liabilities:
Warranty payable 3600
Interest payable 432
Sales tax payable 1500
Notes payable 37000
Total current liabilities 42532

c.

RIPLEY COMPANY
Horizontal Statements Model
Event Assets = Liabilities + Equity Revenue - Expense = Net Income Cash Flow
1 + = + + NA NA - NA = NA + FA
2 + = + + + + - NA = + + OA
3 NA = + + - NA - + = - NA
4 - = - + NA NA - NA = NA - OA
5a. - = - + NA NA - NA = NA - FA
5b. - = NA + - NA - + = - - OA
6 + = + + NA NA - NA = NA + FA
7 - = - + NA NA - NA = NA NA
8 NA = NA + NA NA - NA = NA NA